A 0.314 M solution of a weak base has a pH of 10.70. What is the base hydrolysis constant, K, for the weak base? Kb =
Answer #1

B 0. 3 o31ч - x HBP +04 8 © -tata PH = 1070 - РОН : 19-10 .70 - 3.3 Гон) - 10 33-5-ol x10 К = х2 (6 31ч -7) . (501xto) (o 31ч